OpenStudy (anonymous):

Find an equation for the nth term of a geometric sequence where the second and fifth terms are -24 and 1536, respectively. I know this is basic but can someone explain?

OpenStudy (anonymous):

I know have to use the formula for nth term but I just don't know how to apply it.

OpenStudy (anonymous):

Wait what?

OpenStudy (anonymous):

sequence is \[a,ar, ar^2, ar^3, ar^4, ...\] that is what a geometric sequence looks like

OpenStudy (anonymous):

you are told \(ar=-24,ar^4=1536\)

OpenStudy (anonymous):

Right!

OpenStudy (anonymous):

to find \(r\) divide

OpenStudy (anonymous):

Divide 1539 over -24?

OpenStudy (anonymous):

\[\frac{ar^4}{ar}=\frac{1536}{-24}\]

OpenStudy (anonymous):

So that equals -64.

OpenStudy (anonymous):

so \[\frac{ar^4}{ar}=r^3=-64\]

OpenStudy (anonymous):

solve and get \(r=\sqrt[3]{-64}=-4\)

OpenStudy (anonymous):

So the second term is -24 and the third is -64.

OpenStudy (anonymous):

no you are just finding \(r\)

OpenStudy (anonymous):

I am so lost.. :(

OpenStudy (anonymous):

But I have never seen a cubed root used to find r..

OpenStudy (anonymous):

from algebra, what is \[\frac{ar^4}{ar}\]?

OpenStudy (anonymous):

\[ar^3\]?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

the \(a\)'s cancel as well

OpenStudy (anonymous):

\[r^3\]

OpenStudy (anonymous):

OH. I got it.

OpenStudy (anonymous):

whew !

OpenStudy (anonymous):

So if \[r^3\] is equal to -64

OpenStudy (anonymous):

I'm sorry reading math on here is difficult.

OpenStudy (anonymous):

yeah you got it

OpenStudy (anonymous):

I'm in an upper level math and reviewing this is so weird for me.

OpenStudy (anonymous):

\[r=-4\]

OpenStudy (anonymous):

right because the cube root is -4.

OpenStudy (anonymous):

So if r = -4, the first term is 6?

OpenStudy (anonymous):

ok now we got \(r\) what is \(a\)?

OpenStudy (anonymous):

I said 6...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So it's \[6 * 4^(n+1)\]

OpenStudy (anonymous):

With the n + 1 raised?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

\[a=6, r=-4\]

OpenStudy (anonymous):

terms are \[6, 6\times (-4), 6\times (-4)^2,6\times (-4)^3,...\]

OpenStudy (anonymous):

Am I right?

OpenStudy (anonymous):

I'm only looking for the equation..

OpenStudy (anonymous):

I don't need the actual terms.